# Split Federated Learning for UAV-Enabled Integrated Sensing, Computation, and Communication

### v1

@@ Abstract
Unmanned aerial vehicles (UAVs) with integrated sensing, computation, and communication (ISCC) capabilities have become key enablers of next-generation wireless networks. Federated edge learning (FEL) leverages UAVs as mobile learning agents to collect data, perform local model updates, and contribute to global model aggregation. However, existing UAV-assisted FEL systems face critical challenges, including excessive computational demands, privacy risks, and inefficient communication, primarily due to the requirement for full-model training on resource-constrained UAVs. To deal with aforementioned challenges, we propose Split Federated Learning for UAV-Enabled ISCC (SFLSCC), a novel framework that integrates split federated learning (SFL) into UAV-assisted FEL. SFLSCC optimally partitions model training between UAVs and edge servers, significantly reducing UAVs’ computational burden while preserving data privacy. We conduct a theoretical analysis of UAV deployment, split point selection, data sensing volume, and client-side aggregation frequency, deriving closed-form upper bounds for the convergence gap. Based on these insights, we conceive a joint optimization problem to minimize the energy consumption required to achieve a target model accuracy. Given the non-convex nature of the problem, we develop a low-complexity algorithm to efficiently determine UAV deployment, split point selection, and communication frequency. Extensive simulations on a target motion recognition task validate the effectiveness of SFLSCC, demonstrating superior convergence performance and energy efficiency compared to baseline methods.
@@
Federated edge learning, unmanned aerial vehicle (UAV), split federated learning, integrated sensing, computation, and communication (ISCC), energy-efficient learning.
@@ I Introduction
Unmanned aerial vehicles (UAVs) have evolved beyond their traditional roles, becoming integral to next-generation wireless networks due to their integrated sensing, computation, and communication (ISCC) capabilities [1]. Their mobility and onboard processing capabilities allow them to not only sense and interact with the environment but also support distributed intelligence at the network edge [2]. However, training machine learning (ML) models in such decentralized and dynamic environments presents significant challenges, particularly in terms of data privacy and communication efficiency [3].
Federated edge learning (FEL), as an efficient distributed learning paradigm that enables collaborative model training across edge devices without requiring raw data exchange [4]. UAVs, with their ability to collect real-time data and perform local computation, can play a crucial role in FEL by serving as mobile learning agents [5]. Specifically, they can acquire task-relevant data through onboard sensing, perform local training, and iteratively exchange model updates with a central server, refining a global model in a collaborative and efficient manner. This synergy between UAV-enabled data collection and distributed learning paves the way for more adaptive and intelligent wireless networks.
Despite the promising application prospects of UAV-assisted FEL, practical deployment remains challenging due to several key obstacles. First, UAVs are inherently constrained by limited computational power and battery capacity, making it impractical for them to train full ML models with high-dimensional parameters—such as those in classical models like AlexNet [6] and VGG16 [7]. The high computational demands and energy consumption associated with full-model training often exceed the operational capabilities of resource-limited UAVs, significantly restricting the feasibility of existing FL approaches [8]. Second, in federated learning (FL), the complete model is typically accessible to all involved parties, including UAVs and servers, raising concerns regarding privacy leakage [9]. Additionally, UAV’s deployment affects both data sensing and communication efficiency, which in turn impacts the overall performance of FL systems [5].
@@ I-A Related Works
In this context, the integration of SFL into UAV-assisted FEL systems presents a promising solution. Although no existing work has explored the integration of SFL into UAV-assisted FEL systems, recent studies have investigated fundamental SFL principles, key phenomena, and optimization strategies. A crucial factor is the selection of the split point, which significantly influences the computational load on the client and server, communication overhead, and model convergence speed. Previous research has examined the impact of split point selection. For instance, in [12, 13], the authors developed regression models based on the delay and energy consumption of different split points to determine the optimal division. Wu et al. [14] considered device heterogeneity and network dynamics, optimizing split point selection to minimize training costs. Lin et al. [15] addressed the adaptive selection of split points in multi-device scenarios. Furthermore, the frequency of client aggregation plays a critical role in model convergence and communication overhead. Typically, a higher aggregation frequency accelerates model convergence but also increases communication costs [16]. In [15] and [17], the authors explored the joint optimization of split point selection and aggregation frequency for balancing training accuracy and communication costs. Some studies have also analyzed the theoretical relationship between communication rounds and convergence speed, aiming to minimize overall training costs (in terms of delay or energy consumption) while satisfying desired convergence performance constraints [18, 19, 20].
Inspired by these studies, this work aims to introduce the concept of split federated learning into UAV-Enabled ISCC, taking UAV-driven collaborative intelligence to a new level.
@@ I-B Contributions
This paper proposes a novel split federated learning framework for UAV-enabled integrated sensing, computation, and communication (SFLSCC). The main contributions are as follows:
  • To the best of our knowledge, this is the first work to incorporate SFL into UAV-assisted ISCC system. We theoretically analyze the impact of UAV deployment, split point selection, data sensing volume, and client-side aggregation frequency on the convergence of SFLSCC and derive closed-form upper bounds for the convergence gap.
  • Building upon the convergence analysis, we formulate a joint optimization problem to minimize the energy consumption of the UAV-assisted ISCC system while ensuring a target model accuracy. The optimization jointly considers split point selection, aggregation frequency, UAV deployment strategy, and data sensing volume.
  • Given the non-convex nature of the formulated problem, we design a low-complexity algorithm to efficiently obtain a near-optimal solution. Furthermore, we conduct extensive experiments on a target motion recognition task using a high-fidelity wireless perception simulator. The results demonstrate that SFLSCC achieves superior convergence performance and energy efficiency compared to baseline methods.
The rest of this paper is organized as follows. Section II introduces the system model, while section III presents the convergence analysis of SFLSCC. Section IV formulates the optimization problem and provides the solution algorithm. Subsequently, the simulation results are presented in Section V, and the conclusion is provided in Section VI.
@@ II System Model
The proposed SFLSCC architecture is shown in Fig. 1, comprising an edge server, and a set of $M$ UAVs denoted as =$\mathcal{M}=\{1,2,\ldots,m,\ldots,M\}$, is designed to collaboratively train a ML model for a specific sensing recognition task. Each UAV is equipped with a single-antenna ISAC transceiver that can switch between sensing and communication modes in a time-division manner depending on the mission phase. Specifically, during the sensing phase of round $t$, UAV $m$ senses the target by transmitting a frequency-modulated continuous wave (FMCW) and collect training data, which will be transmitted in the communication phase. However, considering the limits of sensing and communication resources, the full model $\boldsymbol{w}$ which has $L$-layers is divided into the client-side model $\boldsymbol{w}_{c,m}$ and the server-side model $\boldsymbol{w}_{s,m}$. The client-side model reserves neurons of layer $\{1,2,\dots,L_{c}\}$, while the server-side model reserves neurons of layer $\{L_{c}+1,L_{c}+2,\dots,L\}$, where $L_{c}$ represents the split layer. Meanwhile, let $\boldsymbol{u}_{m}=\left(x_{m},y_{m},H\right)$ and $\boldsymbol{u}_{s}=\left(x_{s},y_{s},z_{s}\right)$ denote the three-dimensional coordinates of UAV $m$ and the edge server, respectively.
@@ II-A Split Federated Learning Model
Figure 1: The architecture of SFLSCC.
@@ II-A5 Aggregation of client-side
The UAVs send the client-side models $\boldsymbol{w}_{c,m,t+1}$ to the server every $I$ rounds for model aggregation, which can be denoted as
@@ II-C Target Sensing Model
II-D Energy Consumption Model
Given that the server has abundant and constant energy supply, we only consider the UAVs’ energy consumption, which mainly generates from the process of local training and local data uploading. At the $t$-th round, the energy consumption of UAV $m$ can be given as follows [13].
II-D1 Energy consumption of target sensing
The energy consumed by UAV $m$ for target sensing to collect data can be denoted as
$E_{m,s}=p_{s}T_{m,s}.$ (17)
II-D2 Energy consumption of local training
@@
$T_{m,tr}=\frac{bL_{c}(\xi_{f}+\xi_{b})}{Lf_{m}\varpi}$ (18)
where $f_{m}$ signifies the computing capability (i.e., CPU frequency) of UAV $m$, and $\varpi$ denotes the computing intensity. Therefore, the energy consumption of local training for UAV $m$ can be calculated by
$E_{m,tr}=p_{m,tr}T_{m,tr}$ (19)
where $p_{m,tr}$ is the computation power for local training of UAV $m$.
II-D3 Energy consumption of communication
@@
Then, the energy consumption of smashed data uploading for UAV $m$ is denoted as
$E_{m,com}^{sma}=p_{c}T_{m,com}^{sma}.$ (21)
Meanwhile, the transmission latency to upload local model parameters $|\boldsymbol{w}_{c,m}|$ and the energy consumption can be represented by
$T_{m,com}^{par}=\frac{\boldsymbol{d}_{m}\left(L_{c}\right)}{R_{m}\left(\boldsymbol{u}_{m}\right)},$ (22)
and
$E_{m,com}^{par}=p_{c}T_{m,com}^{par},$ (23)
respectively. And $\boldsymbol{d}_{m}\left(L_{c}\right)=\boldsymbol{A}\left(\frac{L_{c}}{L}\right)^{2}m_{\boldsymbol{w}}$, where $\boldsymbol{A}$ is a positive constant and $m_{\boldsymbol{w}}$ denotes the bits of overall model parameters.
II-D4 Overall energy consumption of $I$ rounds
We consider each round has $e$ epochs, thereby the energy consumption of UAV $m$ can be given by
$E_{m,t}^{epo}=\alpha_{m}^{t}e\left(E_{m,tr}+E_{m,com}^{sma}+E_{m,s}\right).$ (24)
Besides, given that the split training is carried out every round while the local model parameters uploading occurs every $I$ rounds, thus the overall expected energy consumption of UAV $m$ for $I$ rounds can be calculated by
$E_{m}\left(I,L_{c},\boldsymbol{u}_{m},b\right)=\mathbb{E}[IE_{m,t}^{epo}+E_{m,com}^{par}],$ (25)
@@ III Convergence Analysis
In this section, we evaluate the impact of split point selection, aggregation frequency, data sensing volume and locations of UAVs on the proposed SFLSCC through conducting a detailed convergence analysis.
@@ III-B Convergence Analysis
$+32\beta^{2}\eta^{4}I^{2}\sum\limits_{l=1}^{L_{c}}G_{l}^{2}+8\eta^{2}\mathbb{E}\left[\|\nabla\mathcal{L}(\boldsymbol{w}^{t-1})\|^{2}\right].$
@@
We consider the learning rate $\eta$ of proposed SFLSCC satisfies that
@@
$\quad+\frac{2\beta\eta^{2}}{M^{2}}\sum\limits_{m=1}^{M}\Lambda_{m}^{2}+16\beta^{3}\eta^{4}I^{2}\sum\limits_{l=1}^{L_{c}}G_{l}^{2}$
@@
When $\frac{1}{N}\sum\limits_{t=1}^{N}\mathbb{E}\left[\|\nabla\mathcal{L}\left(\mathbf{w}^{t-1}\right)\|^{2}\right]\leq\varepsilon$, it represents the SFLSCC has converged to the target convergence accuracy $\varepsilon$, and the training rounds demanded can be represented as
@@ IV Problem Formulation and Solution Algorithm
As previously highlighted, balancing training performance and energy consumption is critical due to the limited battery capacity of UAVs. To address this, we formulate an energy consumption minimization problem with convergence guarantees by jointly optimizing the aggregation frequency $I$, split point selection $L_{c}$, UAV 3-D coordinates $\{\boldsymbol{u}_{m}\}$, and data sensing volume $b$.
$\mathcal{P}1:$ $\min_{I,L_{c},\{\boldsymbol{u}_{m}\},b}\sum_{m=1}^{M}TE_{m}(I,L_{c},\boldsymbol{u}_{m},b)$ (40a)
@@
where $T=\frac{N}{I}$ represents the total number of communication cycles required to achieve model convergence. Eq. (40b) ensures that the number of training rounds $N$ satisfies the global convergence accuracy, while Eq. (40c) specifies the valid range of the split point $L_{c}$. The constraint Eq. (40d) originates from Corollary 1, which requires that UAVs maintain a consistent target sensing probability. To ensure reliable target sensing quality, we impose the constraint in Eq. (40e). Eqs. (40f) and (40g) ensure that the aggregation frequency $I$ and data sensing volume $b$ are positive integers.
@@
(a) Total energy consumption.
(b) Average test loss.
(c) Average test accuracy.
Figure 3: Comparison of energy consumption and convergence performance across different schemes.
(a) SFLSCC with different $I$.
(b) SFLSCC with different $L_{c}$.
(c) SFLSCC with different $b$.
Figure 4: The impact of different optimization variables on energy consumption of SFLSCC.
(a) SFLSCC with different $I$.
(b) SFLSCC with different $L_{c}$.
(c) SFLSCC with different $b$.
Figure 5: The impact of different optimization variables on the convergence of SFLSCC.
@@
Figure 6: The impact of different environments on energy consumption and convergence of SFLSCC.

### v2

# Split Federated Learning for Low-Altitude Wireless Networks: Joint Sensing, Communication, Computation, and Control Co-design

@@ Abstract
Unmanned aerial vehicles (UAVs) with integrated sensing, communication, computation and control (ISC3) capabilities have become key enablers of next-generation wireless networks. Federated edge learning (FEL) leverages UAVs as mobile learning agents to collect data, perform local model updates, and contribute to global model aggregation. However, existing UAV-assisted FEL systems face critical challenges, including excessive computational demands, privacy risks, and inefficient communication, primarily due to the requirement for full-model training on resource-constrained UAVs. To deal with aforementioned challenges, we propose Split Federated Learning for UAV-Enabled ISC3 (SFLSC3), a novel framework that integrates split federated learning (SFL) into UAV-assisted FEL. SFLSC3 optimally partitions model training between UAVs and edge servers, significantly reducing UAVs’ computational burden while preserving data privacy. We conduct a theoretical analysis of UAV deployment, split point selection, data sensing volume, and client-side aggregation frequency, deriving closed-form upper bounds for the convergence gap. Based on these insights, we conceive a joint optimization problem to minimize the delay required to achieve a target model accuracy. Given the non-convex nature of the problem, we develop a low-complexity algorithm to efficiently determine UAV deployment, split point selection, and communication frequency. Extensive simulations on a target motion recognition task validate the effectiveness of SFLSC3, demonstrating superior convergence and delay performance compared to baseline methods.
@@
Federated edge learning, unmanned aerial vehicle (UAV), split federated learning, integrated sensing, communication, computation, and control (ISC3).
@@ I Introduction
The rapid development of the low-altitude economy (LAE) is driving a wide range of emerging low-altitude applications, including aerial logistics, disaster response, and urban sensing. As a core carrier of such applications, unmanned aerial vehicles (UAVs) have evolved beyond their traditional roles by leveraging their inherent mobility and onboard intelligence to support complex low-altitude operations, thereby giving rise to low-altitude intelligent systems characterized by integrated sensing, communication, computation, and control (ISC3) [23]. Their mobility and onboard processing capabilities allow them to not only sense and interact with the environment but also support distributed intelligence at the network edge [27][4]. However, training machine learning (ML) models in such decentralized and dynamic environments presents significant challenges, particularly in terms of data privacy and communication efficiency [21].
Federated edge learning (FEL), as an efficient distributed learning paradigm that enables collaborative model training across edge devices without requiring raw data exchange [16][7]. UAVs, with their ability to collect real-time data and perform local computation, can play a crucial role in FEL by serving as mobile learning agents [17]. Specifically, they can acquire task-relevant data through onboard sensing, perform local training, and iteratively exchange model updates with a central server, refining a global model in a collaborative and efficient manner. This synergy between UAV-enabled data collection and distributed learning paves the way for more adaptive and intelligent wireless networks.
Despite the promising application prospects of UAV-assisted FEL, practical deployment remains challenging due to several key obstacles. First, UAVs are inherently constrained by limited computational power and battery capacity, making it impractical for them to train full ML models with high-dimensional parameters—such as those in classical models like AlexNet [5] and VGG16 [13]. The high computational demands and delay associated with full-model training often exceed the operational capabilities of resource-limited UAVs, significantly restricting the feasibility of existing FL approaches [14]. Second, in federated learning (FL), the complete model is typically accessible to all involved parties, including UAVs and servers, raising concerns regarding privacy leakage [25]. Additionally, UAV’s deployment affects both data sensing and communication efficiency, which in turn impacts the overall performance of FL systems [17].
@@ I-A Related Works
In this context, the integration of SFL into UAV-assisted FEL systems presents a promising solution. Although no existing work has explored the integration of SFL into UAV-assisted FEL systems, recent studies have investigated fundamental SFL principles, key phenomena, and optimization strategies. A crucial factor is the selection of the split point, which significantly influences the computational load on the client and server, communication overhead, and model convergence speed. Previous research has examined the impact of split point selection. For instance, in [24, 15], the authors developed regression models based on the delay and energy consumption of different split points to determine the optimal division. Wu et al. [22] considered device heterogeneity and network dynamics, optimizing split point selection to minimize training costs. Lin et al. [8] addressed the adaptive selection of split points in multi-device scenarios. Hou et al. [3] proposed the SFL-Integrated Sensing, Computation, and Communication (ISCC) framework, integrating model splitting technology into FL-based UAV ISCC systems, and further incorporated semi-supervised learning into the framework to address the issue of sparse labeled data in UAV networks.
Furthermore, the frequency of client aggregation plays a critical role in model convergence and communication overhead. Typically, a higher aggregation frequency accelerates model convergence but also increases communication costs [20]. In [8] and [26], the authors explored the joint optimization of split point selection and aggregation frequency for balancing training accuracy and communication costs. Some studies have also analyzed the theoretical relationship between communication rounds and convergence speed, aiming to minimize overall training costs (in terms of delay or energy consumption) while satisfying desired convergence performance constraints [12, 10, 19].
Inspired by these studies, this work aims to introduce the concept of split federated learning into UAV-Enabled ISC3, taking UAV-driven collaborative intelligence to a new level.
@@ I-B Contributions
This paper proposes a novel split federated learning framework for UAV-enabled integrated sensing, communication, computation, and control (SFLSC3). The main contributions are as follows:
  • To the best of our knowledge, this is the first work to incorporate SFL into UAV-assisted ISC3 system. We theoretically analyze the impact of UAV deployment, split point selection, data sensing volume, and client-side aggregation frequency on the convergence of SFLSC3 and derive closed-form upper bounds for the convergence gap.
  • Building upon the convergence analysis, we formulate a joint optimization problem to minimize the delay of the UAV-assisted ISC3 system while ensuring a target model accuracy. The optimization jointly considers split point selection, aggregation frequency, UAV deployment strategy, and data sensing volume.
  • Given the non-convex nature of the formulated problem, we design a low-complexity algorithm to efficiently obtain a near-optimal solution. Furthermore, we conduct extensive experiments on a target motion recognition task using a high-fidelity wireless perception simulator. The results demonstrate that SFLSC3 achieves superior convergence performance and latency performance compared to baseline methods.
The rest of this paper is organized as follows. Section II introduces the system model, while section III presents the convergence analysis of SFLSC3. Section IV formulates the optimization problem and provides the solution algorithm. Subsequently, the simulation results are presented in Section V, and the conclusion is provided in Section VI.
@@ II System Model
The proposed SFLSC3 architecture is shown in Fig. 1, comprising an edge server, and a set of $M$ UAVs denoted as $\mathcal{M}=\{1,2,\ldots,m,\ldots,M\}$, is designed to collaboratively train a ML model for a specific sensing recognition task. Each UAV is equipped with a single-antenna ISAC transceiver that can switch between sensing and communication modes in a time-division manner depending on the mission phase. Specifically, during the sensing phase of round $t$, UAV $m$ senses the target by transmitting a frequency-modulated continuous wave (FMCW) and collect training data, which will be transmitted in the communication phase. However, considering the limits of sensing and communication resources, the full model $\boldsymbol{w}$ which has $L$-layers is divided into the client-side model $\boldsymbol{w}_{c,m}$ and the server-side model $\boldsymbol{w}_{s,m}$. The client-side model reserves neurons of layer $\{1,2,\dots,L_{c}\}$, while the server-side model reserves neurons of layer $\{L_{c}+1,L_{c}+2,\dots,L\}$, where $L_{c}$ represents the split layer. Meanwhile, let $\boldsymbol{u}_{m}=\left(x_{m},y_{m},H\right)$ and $\boldsymbol{u}_{s}=\left(x_{s},y_{s},z_{s}\right)$ denote the three-dimensional coordinates of UAV $m$ and the edge server, respectively.
@@ II-A Split Federated Learning Model
Figure 1: The architecture of SFLSC3.
@@ II-A5 Aggregation of client-side
The UAVs send the client-side models $\boldsymbol{w}_{c,m}^{t+1}$ to the server every $I$ rounds for model aggregation, which can be denoted as
@@ II-C Target Sensing Model
II-D Delay Model
At the $t$-th round, the delay of UAV $m$ can be given as follows [15].
II-D1 Delay of target sensing
The delay consumed by UAV $m$ for target sensing to collect data can be denoted as $T_{m,s}$.
II-D2 Delay of local training
@@
$T_{m,tr}=\frac{b(\xi_{f,L_{c}}+\xi_{b,L_{c}})}{f_{m}\varpi}$ (17)
where $f_{m}$ signifies the computing capability (i.e., CPU frequency) of UAV $m$, and $\varpi$ denotes the computing intensity.
II-D3 Delay of communication
@@
Meanwhile, the transmission latency to upload local model parameters $|\boldsymbol{w}_{c,m}|$ can be represented by
$T_{m,com}^{par}=\frac{\boldsymbol{d}_{m}\left(L_{c}\right)}{R_{m}\left(\boldsymbol{u}_{m}\right)}.$ (19)
And $\boldsymbol{d}_{m}\left(L_{c}\right)=\boldsymbol{A}\left(\frac{L_{c}}{L}\right)^{2}m_{\boldsymbol{w}}$, where $\boldsymbol{A}$ is a positive constant and $m_{\boldsymbol{w}}$ denotes the bits of overall model parameters.
II-D4 Overall delay of $I$ rounds
We consider each round has $e$ epochs, thereby the delay of UAV $m$ can be given by
$T_{m,t}^{epo}=\alpha_{m}^{t}e\left(T_{m,tr}+T_{m,com}^{sma}+T_{m,s}\right).$ (20)
Besides, given that the split training is carried out every round while the local model parameters uploading occurs every $I$ rounds, thus the overall expected delay of UAV $m$ for $I$ rounds can be calculated by
$T_{m}\left(I,L_{c},\boldsymbol{u}_{m},b\right)=\mathbb{E}[IT_{m,t}^{epo}+T_{m,com}^{par}],$ (21)
@@ III Convergence Analysis
In this section, we evaluate the impact of split point selection, aggregation frequency, data sensing volume and locations of UAVs on the proposed SFLSC3 through conducting a detailed convergence analysis.
@@ III-B Convergence Analysis
[…]
@@
We consider the learning rate $\eta$ of proposed SFLSC3 satisfies that
@@
[…]
@@
When $\frac{1}{N}\sum\limits_{t=1}^{N}\mathbb{E}\left[\|\nabla\mathcal{L}\left(\mathbf{w}^{t-1}\right)\|^{2}\right]\leq\varepsilon$, it represents the SFLSC3 has converged to the target convergence accuracy $\varepsilon$, and the training rounds demanded can be represented as
@@ IV Problem Formulation and Solution Algorithm
As previously highlighted, balancing training performance and delay is critical due to the limited battery capacity of UAVs. To address this, we formulate an delay minimization problem with convergence guarantees by jointly optimizing the aggregation frequency $I$, split point selection $L_{c}$, UAV 3-D coordinates $\{\boldsymbol{u}_{m}\}$, and data sensing volume $b$.
$\mathcal{P}1:$ $\min_{I,L_{c},\{\boldsymbol{u}_{m}\},b}\sum_{m=1}^{M}\mathcal{E}T_{m}(I,L_{c},\boldsymbol{u}_{m},b)$ (36a)
@@
where $\mathcal{E}=\frac{N}{I}$ represents the total number of communication cycles required to achieve model convergence. Eq. (36b) ensures that the number of training rounds $N$ satisfies the global convergence accuracy, while Eq. (36c) specifies the valid range of the split point $L_{c}$. The constraint Eq. (36d) originates from Corollary 1, which requires that UAVs maintain a consistent target sensing probability. To ensure reliable target sensing quality, we impose the constraint in Eq. (36e). Eqs. (36f) and (36g) ensure that the aggregation frequency $I$ and data sensing volume $b$ are positive integers.
@@
(a) Total Delay.
(b) Average test accuracy vs round.
(c) Average test accuracy vs time.
Figure 3: Comparison of delay and convergence performance across different schemes.
(a) SFLSC3 with different $I$.
(b) SFLSC3 with different $L_{c}$.
(c) SFLSC3 with different $b$.
(d) SFLSC3 with different $\theta_{s}$.
Figure 4: The impact of different optimization variables on delay of SFLSC3.
(a) SFLSC3 with different $I$.
(b) SFLSC3 with different $L_{c}$.
(c) SFLSC3 with different $b$.
(d) SFLSC3 with different $\theta_{s}$.
Figure 5: The impact of different optimization variables on the convergence vs time of SFLSC3.
(a) SFLSC3 with different $I$.
(b) SFLSC3 with different $L_{c}$.
(c) SFLSC3 with different $b$.
(d) SFLSC3 with different $\theta_{s}$.
Figure 6: The impact of different optimization variables on the convergence vs round of SFLSC3.
@@
Figure 7: The impact of different environments on delay and convergence of SFLSC3.